# Query-Efficient Zeroth-Order Algorithms for Nonconvex Optimization

### v1

@@ 4  Zeroth-Order Block Smoothed Gradient Descent Ascent
In this section, we leverage block updates with a variant of GDA, smoothed GDA, to design a new algorithm, and show the best-known convergence result for solving problem (1).
@@ 4.1  Algorithm Design
$K(x,y;z)=f(x,y)+\frac{p}{2}\|x-z\|^{2},$
for some auxiliary variable $z\in\mathbb{R}^{d_{x}}$. The squared term can introduce strong convexity and further smoothness in $x$ with a proper $p$. Then, we will perform gradient descent ascent on the smoothed function $K(x,y;z)$, which is inspired by the first-order smoothed GDA (SGDA) in Zhang et al., (2020). The zeroth-order block smoothed gradient descent ascent (ZOB-SGDA) algorithm is proposed as shown in Algorithm 2. Similarly to ZOB-GDA, we randomly sample a block of dimensions $\mathcal{I}_{k}$ for primal variables and only update the selected dimensions using BCGEs at each step. We denote the partial gradients along the dimensions $\mathcal{I}_{k}$ as
@@
where $\odot$ denotes the Hadamard (element-wise) product. We also denote $G_{x}(x_{k},y_{k};z_{k})=G_{x}^{\mathcal{I}}(x_{k},y_{k};z_{k})$ when $\mathcal{I}=[d_{x}]$. The update of $y_{k}$ follows the same way in ZOB-GDA. Additionally, an extra update for $z_{k}$ is introduced by an averaging step. When $\gamma=1$, it is obvious that ZOB-SGDA can resemble ZOB-GDA.
@@
3:  Randomly sample a set $\mathcal{I}_{k}\subseteq[d_{x}]$ with $|\mathcal{I}_{k}|=b$ and update $x_{k}$ by
@@
4:  Update $y_{k}$ by $y_{k+1}=\mathcal{P}_{\mathcal{Y}}\left[y_{k}+\beta\cdot\nabla_{y}K(x_{k},y_{k};z_{k})\right].$
@@ 4.2  Convergence Analysis of ZOB-SGDA
By properly setting the parameters for ZOB-SGDA, we can establish its convergence result as summarized in Theorem 4.1. Its proof is provided in Appendix B.
@@
Suppose Assumptions 2.1 and 2.3 hold. The sequence $\{(x_{k},y_{k})\}_{k=0}^{K}$ is derived from ZOB-SGDA. Set the parameters $p\geq 3L$, $\sum_{k=0}^{K}r_{k}^{2}\leq\frac{1}{b}$, and $\alpha\leq\frac{1}{p+10L+1}$. Furthermore, let $\beta\leq\min\left\{\frac{1}{12L},\frac{\alpha^{2}(p-L)^{2}}{4L(\sqrt{N}+\alpha(p-L))^{2}}\right\}$, and $\gamma\leq\min\left\{\sqrt{\frac{1}{KN}},\frac{1}{36},\frac{1}{768p\beta}\right\}.$ Then, we have
@@
The results in Theorem 4.1 show that ZOB-SGDA can converge to a stationary point at the convergence of $\mathcal{O}((N/K)^{\frac{1}{4}})$. Note that we do not impose the Lipschitz assumption on $f(x,y)$ in Theorem 4.1. Similarly, based on Theorem 4.1, we have the following corollary to characterize the query complexity of ZOB-SGDA.
@@
We can see from Corollary 4.2 that ZOB-SGDA has the query complexity bound $\mathcal{O}(d_{x}/\epsilon^{4})$, regardless of the choice of block sizes. That means our algorithm can achieve the best-known overall query complexity while maintaining controllable single-step efficiency. For instance, only two queries are required for each step when we set $b=1$, which is more efficient than other CGE-based algorithms that require $\mathcal{O}(d)$ queries for a gradient estimation.
@@ 5  Discussions
Stationary Points of (2) Can Provide Solutions to (1).
Our theoretical results establish convergence guarantees to stationary points of $f(x,y)$ for the proposed algorithms, while the convergence guarantees to the solutions to problem (1) are yet to be established. However, under proper conditions, the stationary points of $f(x,y)$ satisfying $\|\mathfrak{g}(x,y)\|=0$ is also a critical KKT point of problem (1). We provide the following lemma to characterize this property. The definition of critical KKT points and the proof of Lemma 5.1 are provided in Appendix C.
@@
Suppose that $(x,y)\in\mathbb{R}^{d_{x}}\times\mathcal{Y}$ is a stationery point of $f(x,y)$ satisfying $\|\mathfrak{g}(x,y)\|=0$ and $y<\overline{y}$. Then, $x$ is a critical KKT point of problem (1).
The condition $y<\overline{y}$ stems from the gap between our assumption that $\mathcal{Y}$ is bounded and the fact that the multiplier $y\geq 0$ is generally not in practice. This is a common and fundamental gap in the analysis of GDA-type algorithms (Nedić and Ozdaglar,, 2009; Liu et al., 2020b, ; Xu et al.,, 2024), which we believe is an important and interesting future research direction.
Extend Block Updates to Broader Problem Settings.
In our problem (1), we deal with all constraints in the general form $c_{j}(x)\leq 0$. If equality constraints $c_{j}(x)=0$ have to be considered, we can incorporate them by adding two inequalities $c_{j}(x)\geq 0$ and $c_{j}(x)\leq 0$. Besides, we can also consider some simple constraints by constraining the feasible space directly:
@@
and deal with $x\in\mathcal{X}$ by projection, i.e., $x_{k+1}=\mathcal{P}_{\mathcal{X}}[x_{k}-\alpha G_{x}^{\mathcal{I}_{k}}(x_{k},y_{k})]$. We can get the same theoretical results in our analysis when $\mathcal{X}$ is convex and decomposable, i.e., $\mathcal{X}=\prod_{i\in[d_{x}]}\mathcal{X}_{i}$. This requirement originates from the fundamental limit of block/coordinate updates (Lian et al.,, 2016; Jin et al.,, 2023). Note that we need to make the modifications: $\mathfrak{g}_{x}(x,y)=\frac{1}{\alpha}\left(x-\mathcal{P}_{\mathcal{X}}[x-\alpha\nabla_{x}f(x,y)]\right)$, and $\Phi_{1/2L}(x)=\min_{u\in\mathcal{X}}\left\{\Phi(u)+\frac{1}{2\lambda}\|u-x\|^{2}\right\}$ for the stationarity measure. The extended theoretical results are straightforward to establish based on our analysis and the non-expansiveness of projection operators; thus, we omit the detailed analysis in this study.
Our algorithms can also be applied to stochastic cases, i.e., $h(x)=\mathbb{E}[h(x;\xi)],c_{j}(x)=\mathbb{E}[c_{j}(x;\xi)]$, where $\xi$ is a random variable defined in a probability space. However, we cannot expect better convergence guarantees than the RGE-based algorithms, because extra variance arises in BCGEs due to the stochasticity of $\xi$, which diminishes the advantage of controllable variance of our methods.
@@ 6  Numerical Simulations
We validate our algorithms through numerical experiments on an energy management problem in a 141-bus distribution network with $d_{x}=168$ (Khodr et al.,, 2008; Zhou et al.,, 2025). In this problem, the goal is to adjust the load of multiple users within a distribution network to curtail a specific amount of load while minimizing the cost of participating in load curtailment. The detailed problem formulation and experimental settings are provided in Appendix D.
@@
Figure 1: Performance of ZOB-GDA and ZOB-SGDA. (a), (b), and (c) present the objective function value, constraint violation, and stationarity measure of ZOB-GDA. (d), (e), and (f) present the corresponding results for ZOB-SGDA.
@@
First, we apply ZOB-GDA and ZOB-SGDA to solve the problem using different block sizes. Their performance is averaged over 50 repeated runs with different initial parameters and shown in Figure 1. The dark curves represent the average performance, and the shaded areas represent the standard deviation. The results show that both ZOB-GDA and ZOB-SGDA can converge to the same objective function value with the constraint satisfied. Their stationarity measures can both converge to 0, which validates our theoretical guarantees. While different block sizes lead to convergence to the same objective, properly selecting the block sizes may improve the query complexity.
@@
Figure 2: Performance comparison of different algorithms.
We also compare our algorithms (block size $b=10$) with three others, i.e., ZO-MinMax (Liu et al., 2020b, ), SZO-ConEx (Nguyen and Balasubramanian,, 2023), and ZOAGP (Xu et al.,, 2024), which can be applied to solve problem (1). Each algorithm is tested with 50 runs, and the average performance is presented in Figure 2. The results show that our algorithms can both converge to a solution significantly faster than other algorithms. To better compare their query complexities, we present the average number of queries required to generate solutions with different qualities, as in Table 2. It is shown that ZOB-GDA and ZOB-SGDA exhibit highly similar performance under different block sizes, while the complexity bound of the latter is theoretically tighter. Notably, even though different block sizes share the same query complexity bound, proper block sizes may lead to much improved performance (over ten times better than existing methods). Moreover, SZO-ConEx and ZOMinMax have the worst performance due to the large variance of RGEs.
In practice, observations from real systems are often noisy, which can influence the accuracy of gradient estimations. We further test our algorithms under noisy cases to validate their robustness. The test results are provided in Appendix D.3.
@@
Table 2: Average numbers of iterations and queries required to generate solutions with certain levels of relative errors and zero constraint violation. “NaN” means no runs can achieve such a solution.
Relative error 10% 1% 0.1%
Iteration Complexity Iteration Complexity Iteration Complexity
ZOBGDA $b=1$ 722.06 1444.12 1213.48 2426.96 1584.84 3169.68
$b=10$ 73.66 810.26 130.70 1437.70 163.78 1801.58
$b=50$ 21.92 1117.92 57.24 2919.24 76.16 3884.16
$b=168$ 51.48 8700.12 185.92 31420.48 265.71 44905.71
@@
ZOAGP 51.48 8700.12 185.92 31420.48 265.71 44905.71

### v2

@@ 4  Zeroth-Order Block Smoothed Gradient Descent Ascent
In this section, we leverage block updates with a variant of GDA, smoothed GDA, to design a new algorithm, and show its convergence result.
@@ 4.1  Algorithm Design
$H(x,y;z)=f(x,y)+\frac{p}{2}\|x-z\|^{2},$
for some auxiliary variable $z\in\mathbb{R}^{d_{x}}$. The squared term can introduce strong convexity and further smoothness in $x$ with a proper $p$. Then, we will perform GDA on the smoothed function $H(x,y;z)$, which is inspired by the first-order smoothed GDA (SGDA) in Zhang et al. (2020). The zeroth-order block smoothed gradient descent ascent (ZOB-SGDA) algorithm is proposed as shown in Algorithm 2. Similarly to ZOB-GDA, we uniformly sample a block of dimensions $\mathcal{I}_{k}$ from $d_{x}]$ for primal variables and only update the selected dimensions using BCGEs at each step. We denote the partial gradients along the dimensions $\mathcal{I}_{k}$ as
@@
where $\odot$ denotes the Hadamard (element-wise) product. We also denote $G_{x}(x_{k},y_{k};z_{k})=G_{x}^{\mathcal{I}}(x_{k},y_{k};z_{k})$ when $\mathcal{I}=[d_{x}]$. The update of $y_{k}$ follows the same way in ZOB-GDA. Additionally, an extra update for $z_{k}$ is introduced by an averaging step to stabilize the primal dynamics, which is crucial for improving ZOB-GDA. When $\gamma=1$, it is obvious that ZOB-SGDA reduces to ZOB-GDA.
@@
3:  Uniformly sample $\mathcal{I}_{k}\subseteq[d_{x}]$ with $|\mathcal{I}_{k}|=b$ and update $x_{k}$ by
@@
4:  Update $y_{k}$ by $y_{k+1}=\mathcal{P}_{\mathcal{Y}}\left[y_{k}+\beta\cdot\nabla_{y}H(x_{k},y_{k};z_{k})\right].$
@@ 4.2  Convergence Analysis of ZOB-SGDA
By properly setting the parameters for ZOB-SGDA, we can establish its convergence result in the sense of $\|\mathfrak{g}(x_{k},y_{k})\|$ as summarized in Theorem 4.1.
@@
Suppose Assumptions 2.1 and 2.3 hold. The sequence $\{(x_{k},y_{k})\}_{k=0}^{K}$ is derived from ZOB-SGDA. Set the parameters $p\geq 3L$, $\sum_{k=0}^{K}r_{k}^{2}\leq\frac{1}{b}$, and $\alpha\leq\frac{1}{p+10L+1}$. Furthermore, let $\beta\leq\min\big\{\frac{1}{12L},\frac{\alpha^{2}(p-L)^{2}}{4L(\sqrt{N}+\alpha(p-L))^{2}}\big\}$, and $\gamma\leq\min\big\{\sqrt{\frac{1}{KN}},\frac{1}{36},\frac{1}{768p\beta}\big\}.$ Then, we have
@@
The results in Theorem 4.1 show that ZOB-SGDA can converge to a stationary point at the rate of $\mathcal{O}\big(\big(\frac{N}{K}\big)^{\frac{1}{4}}\big)$, which corresponds to a KKT point under proper conditions (see Section 5 for further discussion). Note that Theorem 4.1 does not require the Lipschitz continuity assumption for $f(x,y)$. We provide the proof of Theorem 4.1 in Appendix B.
Similarly, based on Theorem 4.1, we have the following corollary to characterize the query complexity of ZOB-SGDA.
@@
We can see from Corollary 4.2 that ZOB-SGDA has the best-known query complexity bound $\mathcal{O}(d_{x}/\epsilon^{4})$ (Xu et al., 2024), regardless of the choice of block sizes. To the best of our knowledge, this result also matches the best-known complexity bound for first-order single-loop algorithms, up to a factor of $d_{x}$ (Xu et al., 2023; Zhang et al., 2020). In other words, ZOB-SGDA can achieve the best-known overall query complexity bound while maintaining controllable single-step efficiency. Specifically, only two queries are required for each step when we set $b=1$, which is more efficient than CGE-based algorithms that require $\mathcal{O}(d)$ queries for a gradient estimation (Hu and Fu, 2025).
@@ 5  Discussions
In this section, we present the following discussions regarding our algorithms and theoretical results.
First, block size plays an important role in determining the performance of our algorithms. Our theory shows that different block sizes share the same overall query complexity bound. Notably, block size directly governs the trade-off between single-step efficiency and iteration complexity. Smaller blocks reduce the number of queries per iteration but typically require more iterations to achieve certain errors. However, as shown in our following numerical results, properly tuned block sizes can substantially reduce the total number of queries in practice. One may therefore treat $b$ as a hyperparameter and use a few short, early-stopping runs to select a small but effective block size that balances per-iteration cost and overall query complexity. Developing finer-grained complexity bounds that make the dependence on $b$ explicit is an interesting direction for future work.
Second, under suitable conditions, the stationary points of (2) can provide solutions to (1). Our theoretical results establish convergence guarantees to stationary points of $f(x,y)$ for the proposed algorithms, while our goal is to establish the convergence guarantees to the solutions to problem (1). Therefore, we define the KKT gap at a point $(x,y)\in\mathbb{R}^{d_{x}}\times\mathcal{Y}$ as
$\mathcal{K}(x,y)\;:=\;\left\|\nabla_{x}f(x,y)\right\|\;+\;\max_{j\in\mathcal{J}}[c_{j}(x)]_{+}\;+\;\max_{j\in\mathcal{J}}(y)_{j}|c_{j}(x)|,$ (7)
where $\cdot]_{+}:=\max(\cdot,0)$. The first term in (7) matches the primal component of $\|\mathfrak{g}(x,y)\|$ and measures the stationarity; the latter two terms quantify primal infeasibility and complementarity gap. It is obvious that $(x,y)$ is a KKT point of problem (1) when $\mathcal{K}(x,y)=0$. We provide the following lemma to characterize the relationship between $\|\mathfrak{g}(x,y)\|$ and the KKT gap, whose proof is provided in Appendix C.
@@
Suppose that $(x,y)\in\mathbb{R}^{d_{x}}\times\mathcal{Y}$ is any point satisfying $y<\overline{y}$ component-wise. Then, under Assumption 2.2, there exists some constant $C>0$ such that $\mathcal{K}(x,y)\leq C\|\mathfrak{g}(x,y)\|$.
Lemma 5.1 implies that the KKT gaps of our algorithms share the same convergence rate as the stationarity measure $\|\mathfrak{g}(x,y)\|$ under mild conditions. The condition $y<\overline{y}$, in fact, requires that the iterates strictly stay below the upper bound of $\mathcal{Y}$. This is a common and fundamental limitation in the analysis of GDA-type algorithms (Nedić and Ozdaglar, 2009; Liu et al., 2020b; Xu et al., 2023), which we believe is an open problem to be addressed in future work.
Moreover, we can extend block updates to broader problem settings. In our problem (1), we deal with all constraints in the general form $c_{j}(x)\leq 0$. If equality constraints $c_{j}(x)=0$ have to be considered, we can incorporate them by adding two inequalities $c_{j}(x)\leq 0$ and $c_{j}(x)\leq 0$. Besides, we can also consider some simple constraints by constraining the feasible space directly:
@@
and deal with $x\in\mathcal{X}$ by projection, i.e., $x_{k+1}=\mathcal{P}_{\mathcal{X}}[x_{k}-\alpha G_{x}^{\mathcal{I}_{k}}(x_{k},y_{k})]$. We can establish the same theoretical results in our analysis when $\mathcal{X}$ is convex and decomposable, i.e., $\mathcal{X}=\prod_{i\in[d_{x}]}\mathcal{X}_{i}$. This requirement originates from the fundamental limit of block/coordinate updates (Lian et al., 2016; Jin et al., 2026). Note that we need to make the modifications: $\mathfrak{g}_{x}(x,y)=\frac{1}{\alpha}\left(x-\mathcal{P}_{\mathcal{X}}[x-\alpha\nabla_{x}f(x,y)]\right)$, and $\Phi_{1/2L}(x)=\min_{u\in\mathcal{X}}\left\{\Phi(u)+L\|u-x\|^{2}\right\}$ for the stationarity measure. The extended theoretical results are straightforward to establish based on our analysis and the non-expansiveness of projection operators; thus, we omit the detailed analysis in this study.
Our methods can also be applied to stochastic settings, where $h(x)=\mathbb{E}[h(x;\xi)]$ and $c_{j}(x)=\mathbb{E}[c_{j}(x;\xi)]$, i.e., $\xi$ is a random variable defined on a probability space. However, we should not expect stronger convergence guarantees than those achieved by RGE-based algorithms, since the randomness of $\xi$ introduces additional variance into BCGEs, which weakens the benefit of low variance offered by our methods.
@@ 6  Numerical Simulations
In this section, we validate our algorithms through numerical experiments on an energy management problem and a high-dimensional parameter optimization problem. The detailed problem formulations and experimental settings are provided in Appendix D.
6.1  Energy Management Problem
@@
We consider an energy management problem in a 141-bus distribution network with $d_{x}=168$ (Khodr et al., 2008; Zhou et al., 2025). In this problem, the goal is to adjust the load of multiple users within a distribution network to curtail a specific amount of load while minimizing the cost of load adjustment. The problem is formulated as
$\min_{x\in\mathcal{X}}h(x)\qquad\text{s.t.}\;p_{c}(x)\leq D,$
where $x\in\mathbb{R}^{168}$ includes the active and reactive power loads of all buses. $h(x)$ is the objective function concerning operation cost and system safety; $p_{c}(x)$ is the total power exchange with the main grid after curtailment; $D$ is the target power level. Both the objective and constraint functions can only be computed via observations without analytical models.
@@
Figure 1: Performance of ZOB-GDA and ZOB-SGDA in the energy management problem. (a), (b), and (c) present the objective function value, constraint violation, and stationarity measure of ZOB-GDA. (d), (e), and (f) present corresponding results for ZOB-SGDA
@@
First, we apply ZOB-GDA and ZOB-SGDA to solve the problem using different block sizes. Their performance is averaged over 50 repeated runs with different initial points and is shown in Figure 1. The dark curves represent the average performance, and the shaded areas represent the standard deviation. The results show that both ZOB-GDA and ZOB-SGDA can converge to the same objective function value with the constraint satisfied. Their stationarity measures can both converge to 0, which validates our theoretical guarantees. While different block sizes lead to convergence to the same objective function value, properly selecting the block sizes may substantially reduce the total number of queries required. Besides, we test ZOB-SGDA with/without parallel techniques (10 workers) for $b$ directional estimations to compare the wall-clock time to achieve different levels of relative errors (REs) and zero constraint violation (CV), as shown in Table 2. The results indicate that parallelism only provides limited speedup when $b$ is large and, in many cases, even leads to worse wall-clock performance for ZOB-SGDA. It may stem from the fact that each function evaluation is inexpensive, and the additional time for parallel processing offsets the saved time due to parallel computing.
Table 2: Computational Time of ZOB-SGDA under Parallel and Non-parallel Settings (Unit: seconds)
$b=1$ $b=10$ $b=50$ $b=168$
RE Parallel Non-parallel Parallel Non-parallel Parallel Non-parallel Parallel Non-parallel
10% 19.38 6.99 4.19 1.12 3.34 0.71 4.84 4.95
1% 31.88 11.82 5.87 1.75 6.06 1.94 17.04 17.28
0.1% 42.57 15.89 6.93 2.22 7.43 2.66 22.34 24.92
@@
Figure 2: Performance comparison of different algorithms in the energy management problem
@@
Table 3: Average numbers of iterations and queries required to generate certain solutions in the energy management problem (“NaN” means no runs can achieve such a solution)
RE 10% 1% 0.1%
Iteration Queries Iteration Queries Iteration Queries
ZOB-GDA $b=1$ 722.06 1444.12 1213.48 2426.96 1584.84 3169.68
$b=10$ 75.00 825.00 138.08 1518.88 182.00 2002.00
$b=50$ 24.32 1240.32 68.02 3469.02 92.24 4704.24
$b=168$ 60.06 10150.14 216.68 36618.92 285.90 48317.10
@@
ZOAGP 45.14 7628.66 162.82 27516.58 255.96 43257.24
We also compare our algorithms (block size $b=10$) with three others, i.e., ZO-MinMax (Liu et al., 2020b), SZO-ConEx (Nguyen and Balasubramanian, 2023), and ZOAGP (Xu et al., 2024), which can be applied to solve problem (1). Each algorithm is tested with 50 runs, and the average performance is presented in Figure 2. The results show that our algorithms can both converge to a solution significantly faster than others. To better compare their query complexities, we present the average number of queries required to generate solutions with zero CV and different REs, as shown in Table 3. ZOB-GDA and ZOB-SGDA exhibit highly similar performance under different block sizes, while the complexity bound of the latter is theoretically tighter. Notably, proper block sizes may lead to much improved query complexities (about 10 times better than existing methods). Moreover, SZO-ConEx and ZO-MinMax have much worse performance due to the large variance of RGEs.
Based on the numerical results, we highlight two key observations. First, although all choices of block size $b$ share the same query complexity bound, their empirical convergence behaviors can differ substantially. This is because our query complexity bounds provide only worst-case guarantees, whereas in practice, the algorithms may perform significantly better with an appropriately chosen $b$. We recommend starting with a small block size and tuning by early stopping runs until a good empirical trade-off between single-step cost and convergence speed is observed. Second, ZOB-GDA exhibits performance comparable to that of ZOB-SGDA in our experiments, despite the fact that ZOB-SGDA enjoys a strictly better query complexity bound. This observation again suggests that the theoretical results are conservative and may not fully capture the typical behavior encountered in practice. The situation is similar to the first-order setting, where GDA is theoretically dominated by its variants, such as extra-gradient methods, yet often performs competitively and can enjoy strong empirical performance except in certain pathological cases.
6.2  High-Dimensional Parameter Optimization
We further test our algorithms in a 1000-dimensional parameter optimization problem to validate the capability of our methods for scalable problems. We consider the following parameter optimization problem:
$\min_{x\in\mathbb{R}^{1000}}$ $h(x)=0.5\cdot\|Bx\|^{2}+0.1\cdot\sum_{i}x^{4}(i)$
s.t. $c(x)=(1+e^{-q^{T}x})^{-1}-\bar{c}\leq 0,$
where $B\in\mathbb{R}^{1000\times 1000}$ is a matrix, $q\in\mathbb{R}^{1000}$ is a vector, and $\bar{c}$ is a scalar. The detailed parameter settings for the problem and algorithms are provided in Appendix D. By the early-stopping rule, we set the block size $b=30$ for our algorithms, which yields good performance.
Figure 3: Performance comparison of different algorithms in the parameter optimization problem
Table 4: Average numbers of queries required to generate certain solutions in the parameter optimization problem (“NaN” means no runs can achieve such a solution)
RE / CV 10% / 0.1 1% / 0.01 0.1% / 0.001
Iteration Queries Iteration Queries Iteration Queries
ZOB-GDA ($b=30$) 1853.00 57443.00 4081.70 126532.70 6321.30 195960.30
ZOB-SGDA ($b=30$) 1704.10 52827.10 3795.55 117662.05 5876.90 182183.90
ZO-MinMax 125127.85 250255.70 279401.10 558802.20 436177.55 872355.10
SZO-ConEx 172631.15 690525.60 NaN NaN NaN NaN
ZOAGP 234.00 234234.00 556.45 557006.45 887.30 888187.30
The results are summarized in Figure 3 and Table 4. Except for SZO-ConEx, all algorithms can generate solutions that meet the prescribed levels of RE and CV, whereas our algorithms converge substantially faster in terms of query efficiency. Notably, the performance of ZOB-SGDA is slightly better than ZOB-GDA in number of queries. In contrast, SZO-ConEx fails to generate high-accuracy solutions, and its performance is significantly worse than that of our algorithms. Interestingly, although ZOAGP and ZO-MinMax apply CGEs and RGEs, respectively, they exhibit similar performance. This suggests that the CGE-based algorithm does not necessarily converge faster in high-dimensional cases, even though CGEs typically enjoy lower variance.